zarkam21:

?. A race car has a centripetal acceleration of 18.75 m/s^2 as it goes around a curve. If the curve is a circle with radius 12 m, what is the speed of the car?

zarkam21:

1 attachment
Vocaloid:

a = v^2/r again

zarkam21:

18.75=12^2/r

Vocaloid:

12 is the radius not the velocity

zarkam21:

are we solving for velocity

Vocaloid:

yes

zarkam21:

18.75=v/12

zarkam21:

B

Vocaloid:

good